OpenStudy (anonymous):

sin^4(15) + cos^4(15) =

OpenStudy (anonymous):

sin^4(15)+cos^4(15)=(sin^2(15)+cos^2(15))^2-2*sin^2(15)*cos^2(15)

OpenStudy (anonymous):

7\8

OpenStudy (anonymous):

sin^2(15)+cos^2(15)=1 as sin^2(x)+cos^2(x)=1

OpenStudy (anonymous):

and 2sinx*cosx=sin2x

OpenStudy (anonymous):

could you explain it by steps and which methods did you use

OpenStudy (anonymous):

that one makes me feel discomfort with the question

OpenStudy (anonymous):

sin^4(15)+cos^4(15)=(1)^2-2*sin^2(15)*cos^2(15) =1-(1/2)2*2*sin^2(15)*cos^2(15) =1-(1/2)*(2*sin15*cos15)^2 =1-(1/2)(sin(2*15))^2 =1-(1/2)sin^2(30) =1-(1/2)*(1/2)^2 =1-1/2*1/4 =1-1/8 =7/8

OpenStudy (anonymous):

it is in form of sin^4x+cos^4x

OpenStudy (anonymous):

thank you very much)))

OpenStudy (anonymous):

so try to convert in sin^2x and cos^2x